### v1

# How Fair is Your Diffusion Recommender Model?

Abstract.
Diffusion-based recommender systems have recently proven to outperform traditional generative recommendation approaches, such as variational autoencoders and generative adversarial networks. Nevertheless, the machine learning literature has raised several concerns regarding the possibility that diffusion models, while learning the distribution of data samples, may inadvertently carry information bias and lead to unfair outcomes. In light of this aspect, and considering the relevance that fairness has held in recommendations over the last few decades, we conduct one of the first fairness investigations in the literature on DiffRec, a pioneer approach in diffusion-based recommendation. First, we propose an experimental setting involving DiffRec (and its variant L-DiffRec) along with nine state-of-the-art recommendation models, two popular recommendation datasets from the fairness-aware literature, and six metrics accounting for accuracy and consumer/provider fairness. Then, we perform a twofold analysis, one assessing models’ performance under accuracy and recommendation fairness separately, and the other identifying if and to what extent such metrics can strike a performance trade-off. Experimental results from both studies confirm the initial unfairness warnings but pave the way for how to address them in future research directions.
1. Introduction and motivation
Research in recommender systems (RSs) has recently been shaped by the latest advances in deep learning, e.g., graph neural networks (Wu et al., 2023a), large language models (Wu et al., 2023b), and diffusion models (Wang et al., 2023). While the first two families have already received significant attention in the literature, diffusion-based RSs are still in early stages.
As of today, diffusion models (Ho et al., 2020; Yang et al., 2024; Cao et al., 2024) have settled as the leading solution in generative AI (Goodfellow et al., 2020) and revolutionized the paradigm of generating new (but realistic) data with several applications in computer vision (Rombach et al., 2022; Ho et al., 2020), natural language processing (Austin et al., 2021; Li et al., 2022), graph data (Niu et al., 2020; Liu et al., 2023), chemistry, and bioinformatics (Xu et al., 2022; Hoogeboom et al., 2022). The core principle behind diffusion models lies in their twofold learning procedure, involving a forward process where samples from a real-data distribution are iteratively corrupted by adding noise, e.g., Gaussian noise, and a reverse process where the model is trained to reconstruct the original data by denoising the corrupted samples.
As observed by Wang et al. (2023), the diffusion model framework is well-suited for recommendation settings, where it may effectively learn to denoise users’ implicit feedback containing noisy interactions, i.e., false positives. Their approach, called DiffRec, shows improved performance to traditional generative RSs, e.g., those based on variational autoencoders (Liang et al., 2018) and generative adversarial networks (Wang et al., 2017) by addressing their limited representational power (the former) or unstable training procedures (the latter). Similarly, other works have proposed to leverage diffusion models for specific recommendation tasks such as point-of-interests (Qin et al., 2024), sequential (Li et al., 2023; Ma et al., 2024), or knowledge graph-based recommendation (Jiang et al., 2024).
Diffusion models have achieved widespread success in many domains and applications, but they are a relatively recent trend in machine learning, still facing several open challenges. For instance, some works have raised concerns regarding unfairness that could affect diffusion models. As observed in generative AI, bias is intrinsic in real-world data, and diffusion models might generate unfair outcomes while learning from such a biased data. In their seminal paper, Friedrich et al. (2023) discuss how stable diffusion methods are prone to generate synthetic images that lack diversity, e.g., firefighters are only associated with white males. Similarly, other works, such as Bianchi et al. (2023), discuss how textual prompts with trait annotations may produce stereotypical images, while Cho et al. (2023) observe bias in the generated images in terms of people color tones and occupations. Recent papers have proposed possible countermeasures, such as introducing fair guidance instructions during the data generation process (Friedrich et al., 2023), obfuscating the sensitive attributes (Choi et al., 2024), or adopting an ad-hoc fine-tuning (Shen et al., 2023).
In light of the above, a natural question arises: “Can diffusion models be harmful to recommendation fairness likewise?”. The question is pertinent given the large relevance that bias and fairness have held in recommendations for decades (Anelli et al., 2022; Mansoury et al., 2022; Sun et al., 2020; Zhao et al., 2022b; Boratto et al., 2023b). Thus, we consider it is imperative to rigorously evaluate the performance of popular diffusion-based RSs, with a specific focus on potential unfairness issues, as there is currently no literature on this topic. Indeed, it is worth underlining that such recommendation models are still in their pioneer era. A thorough investigation of their performance can help identify and address observed pitfalls.
To this end, in the current work, we select one of the pioneer approaches in diffusion-based recommendation, DiffRec (Wang et al., 2023) (and its variant L-DiffRec), to ultimately answer the following research question: “Is DiffRec a fair recommender system compared to the state-of-the-art?”. To conduct our investigation, we consider two datasets (with users’ attribute data) and nine additional state-of-the-art recommender systems. Our purpose is to benchmark their performance on six recommendation metrics accounting for accuracy, consumer (Boratto et al., 2022; Anelli et al., 2023; Vassøy and Langseth, 2024), and provider fairness (Boratto et al., 2024; Naghiaei et al., 2022; Mansoury et al., 2022). Concretely, we run a twofold performance analysis, where we first measure each performance metric separately, and then in a trade-off setting. We believe this twofold perspective can help identify performance patterns by encompassing all actors (consumers and providers) involved in the recommendation scenario (Anelli et al., 2023).
Results depict a worrying scenario, outlining that fairness concerns raised in machine learning could also hold for the recommendation domain. Nonetheless, introducing additional components in the diffusion-based process, as for L-DiffRec, seems to mitigate the negative effects of unfairness; indeed, these findings pave the way to novel future directions in diffusion-based recommendation where fairness is preserved. Despite being a preliminary investigation, to the best of our knowledge, this is the first attempt in the literature to provide a rigorous and fairness-aware evaluation perspective on diffusion-based recommendation.
2. Background
This section provides the background notions regarding the two main topics of this paper, namely, diffusion-based recommendation (with a special focus on DiffRec (Wang et al., 2023)) and the evaluation metrics to assess fairness (consumer and provider sides) in recommendation.
2.1. Diffusion recommender model
Similarly to diffusion models, DiffRec (Wang et al., 2023) proposes to treat recommendation as a denoising process of historical user-item interactions, which may be extremely noisy, i.e., contain false positives. In detail, the recommendation model is trained to first corrupt users’ interactions through noise (forward process) and then reconstruct them via an ad-hoc denoising procedure (reverse process). Differently from traditional diffusion models employed on images, users’ historical interactions are not corrupted into pure noise, but the ratio of added noise decreases consistently.
Let and be the sets of users and items in the system, respectively, with as the vector of user-item interactions for user . Each vector entry denotes user interacted with item , otherwise . Then, the forward process, in a Markov chain process, corrupts the user interaction vector into steps by gradually adding Gaussian noise. The formulation for any step is: , where (with a notation abuse) we indicate and set , , at the initial step, while stands for the Gaussian distribution, controls the magnitude of the Gaussian noise added at each step , and is the standard deviation of the distribution. Conversely, the reverse process involves training any machine learning model, e.g., a neural network, to approximate the mean and standard deviation of the data distribution and recover the original user interaction vector. Again, for any step , we have: , where and are the learned mean and standard deviation through the machine learning model having parameters .
Furthermore, to address large-scale recommendation (a pretty common scenario in real-world applications), Wang et al. (2023) also propose L-DiffRec, a lighter-weight version of DiffRec. This variant first clusters pre-trained item embeddings, e.g., from LightGCN, into clusters. Afterward, the initial user interaction vector is split into sub-vectors according to the corresponding -th item cluster. Then, a variational autoencoder is trained to compress each of user interaction sub-vectors , so that the diffusion, i.e., forward/reverse processes, occurs in a low-dimensional latent space, ideally in parallel computation across all splits. Finally, the original user interaction vector is reconstructed through the decoding process of the previous variational autoencoder.
Note that in Wang et al. (2023), the authors eventually propose a second variant to DiffRec, called T-DiffRec, which can encode the temporal information during the training. However, T-DiffRec is specialized for sequential recommendation, which would require a dedicated study to compare its performance with other analogous models. To this end, we exclude T-DiffRec from our experiments, leaving its adoption to future extensions of our study.
Our analysis aims to offer a preliminary outline on the fairness performance of diffusion-based recommender systems. Although other works have recently contributed with novel diffusion-based solutions (Qin et al., 2024; Li et al., 2023; Ma et al., 2024; Jiang et al., 2024), DiffRec could be rightly considered as the pioneer approach in this domain for chronological reasons, In light of this, we selected only this model and plan to extend our investigation to other diffusion models in forthcoming works.
2.2. Fairness evaluation in recommendation
Recent studies focused on assessing fairness under the notion of demographic parity (Boratto et al., 2022; Naghiaei et al., 2022; Vassøy and Langseth, 2024). On the one hand (Boratto et al., 2023a; Naghiaei et al., 2022), we operationalize demographic parity on the consumer side as the absolute difference in recommendation utility between users’ groups. We select the respective fairness counterparts of nDCG and Recall, namely nDCG and Recall. On the other hand, provider fairness is estimated as the absolute difference in exposure between items’ groups (Exp) (Naghiaei et al., 2022; Boratto et al., 2024) and the average percentage of long-tail items in the top- lists (APLT) (Anelli et al., 2023; Mansoury et al., 2022). Lower values of these metrics reflect fairer recommendations, except for APLT. For page constraints, we do not dive deep into their formulations. The interested reader may refer to the above-cited papers for a detailed presentation.
3. Experimental study
In this section, we present our proposed experimental study aimed at assessing fairness in diffusion-based recommendation. First, we describe the baselines and datasets adopted for the analysis. Then, we provide reproducibility details for the conducted experiments.
3.1. Baselines
We select nine state-of-the-art recommendation models along with DiffRec and L-DiffRec: BPRMF (Rendle et al., 2009), ItemkNN (Sarwar et al., 2001), NeuMF (He et al., 2017), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), XSimGCL (Yu et al., 2024), EASE (Steck, 2019), MultiVAE (Liang et al., 2018), and RecVAE (Shenbin et al., 2020). As reported in Table 1, our selection depends on the following criteria: (i) the baselines are well-established solutions, accepted at top-tier conferences and journals; (ii) they cover a diverse set of recommendation families, such as graph neural networks and generative models; (iii) many approaches have been already analyzed under recommendation fairness, highlighting an existing literature gap in diffusion-based recommendation.
3.2. Datasets
We assess the performance of the selected baselines on MovieLens-1M (ML1M) (Harper and Konstan, 2016) and Foursquare Tokyo (FTKY) (Yang et al., 2016), which are frequently adopted in fairness-aware literature in recommendation (Vassøy and Langseth, 2024; Boratto et al., 2023a; Liu et al., 2021). While ML1M is a popular recommendation dataset that collects movies and users’ preferences on them, FTKY involves interactions represented by users’ check-ins at points of interest in Tokyo. In both datasets, users’ sensitive information comes in a binary gender label. Thus, fairness analyses on the consumer side adopt the gender attribute included in both datasets, whereas provider fairness is assessed on a binary partition of the items’ set, namely short-head (popular) and long-tail (niche) items. Following (Boratto et al., 2024; Naghiaei et al., 2022; Mansoury et al., 2022), we select the 20% most popular items as short-head and the rest as long-tail. Dataset statistics are reported in Table 2.
3.3. Reproducibility
Datasets are split into train, validation, and test sets (70%/10%/20%) following the temporal hold-out method. Then, for the models training and evaluation, we rely on the framework RecBole (Zhao et al., 2022a) for training and evaluating all the baselines and, for the two diffusion-based recommender systems, we adopt their original code implementations111https://github.com/YiyanXu/DiffRec.. All recommendation baselines are trained with a grid-search exploration of the hyper-parameters by following the original papers and selecting the hyper-parameter configuration providing the highest value of the Recall@20 on the validation set.
| Models | Year | Venue | Analyzed fairness in |
| BPRMF (Rendle et al., 2009) | 2009 | UAI | (Anelli et al., 2022, 2023; Mansoury et al., 2022; Sun et al., 2020; Zhao et al., 2022b; Boratto et al., 2022) |
| ItemkNN (Sarwar et al., 2001) | 2001 | WWW | (Anelli et al., 2022; Boratto et al., 2023a, b, 2022) |
| NeuMF (He et al., 2017) | 2017 | WWW | (Anelli et al., 2022; Boratto et al., 2023a, b; Mansoury et al., 2022; Sun et al., 2020; Boratto et al., 2022) |
| LightGCN (He et al., 2020) | 2020 | SIGIR | (Anelli et al., 2023; Boratto et al., 2023a; Zhao et al., 2022b) |
| UltraGCN (Mao et al., 2021) | 2021 | CIKM | (Anelli et al., 2023) |
| XSimGCL (Yu et al., 2024) | 2024 | TKDE | ✗ |
| EASE (Steck, 2019) | 2019 | WWW | (Anelli et al., 2022) |
| MultiVAE (Liang et al., 2018) | 2018 | WWW | (Anelli et al., 2022; Togashi et al., 2024) |
| RecVAE (Shenbin et al., 2020) | 2020 | WSDM | ✗ |
| DiffRec (Wang et al., 2023) | 2023 | SIGIR | ✗ |
| L-DiffRec (Wang et al., 2023) | 2023 | SIGIR | ✗ |
| Datasets | ML1M (Harper and Konstan, 2016) | FTKY (Yang et al., 2016) |
| # Users | 6,040 | 7,240 |
| # Items | 3,650 | 5,779 |
| # Interactions | 1,000,130 | 353,769 |
| Sparsity | 95.46% | 99.15% |
| Gender Repr. | M : 71.7%; F : 28.3% | M : 87.9%; F : 12.1% |
| Min. Interactions | 20 | 20 |
4. Results and discussion
In this section, we aim to answer the following research question: “Is DiffRec a fair recommender system compared to the state-of-the-art?”. To this end, we conduct a twofold analysis. First, we investigate the accuracy and fairness performance separately, and then we measure the trade-off among performance metrics to examine whether DiffRec is suitable to provide both fair and useful recommendations.
| Metrics | BPRMF | ItemkNN | NeuMF | LightGCN | UltraGCN | XSimGCL | EASE | MultiVAE | RecVAE | DiffRec | L-DiffRec | |
| ML1M | Recall () | 10.67 | 9.87 | 10.44 | 10.16 | 10.74 | 10.46 | 6.45 | 8.25 | 9.68 | 10.71 | 10.38 |
| nDCG () | 12.94 | 12.22 | 12.67 | 13.09 | 12.82 | 12.80 | 8.62 | 12.04 | 12.86 | 13.19 | 12.85 | |
| Recall () | 0.43 | 1.08 | 0.32 | 0.19 | 0.06 | 0.14 | 0.70 | 0.48 | 0.53 | 0.86 | 0.66 | |
| nDCG () | 1.57 | 1.98 | 1.54 | 1.43 | 1.37 | 1.34 | 1.66 | 2.43 | 2.07 | 2.03 | 2.25 | |
| APLT () | 12.47 | 8.01 | 11.75 | 8.11 | 13.73 | 10.50 | 9.00 | 2.63 | 5.38 | 6.18 | 10.81 | |
| Exp () | 86.29 | 91.18 | 87.36 | 91.32 | 84.98 | 88.78 | 90.54 | 97.02 | 94.10 | 93.67 | 87.26 | |
| FTKY | Recall () | 12.17 | 11.77 | 12.48 | 11.03 | 12.22 | 13.07 | 6.19 | 10.06 | 12.30 | 13.10 | 12.48 |
| nDCG () | 10.73 | 10.48 | 10.93 | 9.87 | 10.85 | 11.37 | 5.48 | 9.21 | 10.90 | 11.53 | 10.77 | |
| Recall () | 0.94 | 0.69 | 0.59 | 0.77 | 1.05 | 0.98 | 0.40 | 0.97 | 1.00 | 0.80 | 0.61 | |
| nDCG () | 0.22 | 0.19 | 0.09 | 0.47 | 0.29 | 0.23 | 0.10 | 0.59 | 0.29 | 0.22 | 0.02 | |
| APLT () | 9.93 | 18.27 | 10.20 | 7.42 | 5.36 | 11.06 | 6.32 | 3.72 | 7.27 | 10.75 | 16.97 | |
| Exp () | 89.27 | 78.91 | 89.39 | 91.87 | 94.32 | 88.25 | 93.56 | 95.75 | 92.03 | 88.57 | 80.85 |
4.1. Single-metric performance
Table 3 shows the results of the selected baselines on ML1M and FTKY in terms of recommendation accuracy, i.e., Recall and nDCG, consumer fairness, i.e., Recall and nDCG, and provider fairness, i.e., APLT and Exp. We recall that all metrics are calculated on top-20 recommendation lists, and are expressed as percentages (%).
On the recommendation accuracy level, the results are almost aligned with the findings from the related literature, where graph- and diffusion-based recommender systems generally dominate over the other baselines on both recommendation datasets. DiffRec shows slightly superior performance than its variant L-DiffRec, but the latter is more computationally lightweight. This gap in performance between the two DiffRec variants might be due to the diffusion process being conducted in the latent space by L-DiffRec. Specifically, the variational autoencoders employed by L-DiffRec compress the users’ preferences signals, which might lead to information loss and a non-perfect reconstruction in the decoding process. On the other hand, DiffRec forward and reverse processes fully capture the discrete preference information, without dealing with the potential loss of details caused by information bottleneck.
Moving to the experiments on fairness, results exhibit distinct trends across the datasets. On ML1M, graph-based approaches such as UltraGCN and XSimGCL are able to provide high-utility and fair recommendations, while the diffusion models settle among the unfairest approaches across all the baselines. On FTKY, the fairest recommender systems are instead the generative ones, such as EASE and L-DiffRec, and the traditional ones, such as ItemKNN and NeuMF. The diverse trends in consumer fairness across ML1M and FTKY might derive from their different domains. Specifically, users’ interactions in the movie domain are driven by intrinsically individual preferences towards specific movie categories, such as those sharing the same actors or the same genre, whereas venue check-ins in FTKY might derive from goings out decided as groups, e.g., with friends or partners. Biases in movies’ preferences might be exacerbated by the RSs compared with venues’ preferences.
The consumer unfairness exhibited by DiffRec on ML1M underscores that diffusion models are prone to increase the bias impact, as also emphasized by other studies in different domains. However, a common observation across the datasets is that the L-DiffRec variant reports fairer recommendations than DiffRec and the other family-related generative models, e.g., MultiVAE and RecVAE. Particularly, regarding APLT and Exp, L-DiffRec exhibits fairer outcomes on the provider side. This might be attributed to the intrinsic clustering operation of L-DiffRec, which shifts the learning process from following the common popularity bias to capturing more nuanced patterns across items’ categories.
4.2. Trade-off analysis
As the three analyzed properties, i.e., accuracy, consumer, and provider fairness, represent different and possibly contrasting sides of the recommendation scenario, we extend our previous analysis by assessing the trade-off among them, as recently done by related studies (Anelli et al., 2023; Boratto et al., 2023b). Indeed, we believe the current analysis may suitably complement the findings from the single-metric study conducted above, motivating and explaining from other complementary perspectives the previous performance trends.
The Kiviat diagrams in Figure 1 represent the performance for both datasets of five representative baselines, i.e., ItemkNN, XSimGCL, RecVAE, DiffRec, and L-DiffRec, on three metrics, i.e., nDCG, nDCG, and APLT, accounting for accuracy, consumer, and provider fairness. To provide an easier interpretation of the plots, we follow (Anelli et al., 2023) and normalize the metrics into [0, 1] through min-max normalization and rescale nDCG to adhere to the principle that “the higher the values the better the performance”.
On the ML1M dataset, most of the findings from the single-metric analysis are generally confirmed. Indeed, we observe that XSimGCL is the most balanced baseline across all models, being capable of providing an optimal trade-off among all metrics. Conversely, ItemkNN, RecVAE, and DiffRec cannot generally increase one metric side without harming the other ones. Interestingly, though, while ItemkNN tends to maximize fairness over the accuracy, RecVAE and DiffRec, i.e., generative and diffusion-based models, do the opposite. This becomes especially evident on DiffRec, where the model is largely improving accuracy at the detriment of provider/consumer fairness. The exception to the trend is L-DiffRec, reaching the second-best trade-off soon after XSimGCL. In fact, on FTKY, L-DiffRec exhibits even the best trade-off compared to the selected baselines.
Thus, the findings from the trade-off analysis, with a more clarifying visualization, confirm the unfairness warnings raised in the single-metric study but, at the same time, suggest there might be room for improvement in diffusion-based recommendation toward more fairness-aware recommendations.
5. Conclusion and future work
This work settles as a preliminary attempt to rigorously assess the performance of diffusion-based recommender systems, specifically focusing on one of the pioneer approaches, DiffRec, and its variant L-DiffRec. The relevance of our study is emphasized by the machine learning literature that has recently observed that the intrinsic bias within the data may cause diffusion models to generate unfair outcomes. To this end, we decided to tailor our investigation to the performance of (L-)DiffRec under common fairness properties in recommender systems. Specifically, we conducted an extensive analysis involving two recommendation datasets (with sensitive information), nine state-of-the-art recommendation systems, and six recommendation metrics accounting for accuracy, consumer, and provider fairness. First, in a single-metrics evaluation, results generally outlined that DiffRec might exhibit unfairness issues; however, such issues are mitigated by the L-DiffRec variant, especially on the provider side. Then, a complementary trade-off analysis among accuracy and fairness properties confirmed that L-DiffRec might be the key to reaching optimal metrics balancing. We believe the take-home message is that diffusion models cannot be plugged into recommender systems as they are, since they would bring (and even exacerbate) the unfairness issues already observed in machine learning. However, with careful and tailored modifications to the rationale and models’ architecture, e.g., L-DiffRec, the issues might be properly addressed. To confirm our final assumption, future works will aim to extend our study to the other DiffRec variant, T-DiffRec, as well as other diffusion-based recommender systems in the literature.
Acknowledgements.
Daniele Malitesta and Fragkiskos D. Malliaros acknowledge the support by ANR (French National Research Agency) under the JCJC project GraphIA (ANR-20-CE23-0009-01). The authors acknowledge Inria for the availability of high-performance computing resources.References
- (1)
- Anelli et al. (2022) Vito Walter Anelli, Alejandro Bellogín, Tommaso Di Noia, Dietmar Jannach, and Claudio Pomo. 2022. Top-N Recommendation Algorithms: A Quest for the State-of-the-Art. In UMAP. ACM, 121–131.
- Anelli et al. (2023) Vito Walter Anelli, Yashar Deldjoo, Tommaso Di Noia, Daniele Malitesta, Vincenzo Paparella, and Claudio Pomo. 2023. Auditing Consumer- and Producer-Fairness in Graph Collaborative Filtering. In ECIR (1) (Lecture Notes in Computer Science, Vol. 13980). Springer, 33–48.
- Austin et al. (2021) Jacob Austin, Daniel D. Johnson, Jonathan Ho, Daniel Tarlow, and Rianne van den Berg. 2021. Structured Denoising Diffusion Models in Discrete State-Spaces. In NeurIPS. 17981–17993.
- Bianchi et al. (2023) Federico Bianchi, Pratyusha Kalluri, Esin Durmus, Faisal Ladhak, Myra Cheng, Debora Nozza, Tatsunori Hashimoto, Dan Jurafsky, James Zou, and Aylin Caliskan. 2023. Easily Accessible Text-to-Image Generation Amplifies Demographic Stereotypes at Large Scale. In FAccT. ACM, 1493–1504.
- Boratto et al. (2023a) Ludovico Boratto, Francesco Fabbri, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2023a. Counterfactual Graph Augmentation for Consumer Unfairness Mitigation in Recommender Systems. In CIKM. ACM, 3753–3757.
- Boratto et al. (2024) Ludovico Boratto, Francesco Fabbri, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2024. Robustness in Fairness Against Edge-Level Perturbations in GNN-Based Recommendation. In ECIR (3) (Lecture Notes in Computer Science, Vol. 14610). Springer, 38–55.
- Boratto et al. (2022) Ludovico Boratto, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2022. Consumer Fairness in Recommender Systems: Contextualizing Definitions and Mitigations. In ECIR (1) (Lecture Notes in Computer Science, Vol. 13185). Springer, 552–566.
- Boratto et al. (2023b) Ludovico Boratto, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2023b. Practical perspectives of consumer fairness in recommendation. Inf. Process. Manag. 60, 2 (2023), 103208.
- Cao et al. (2024) Hanqun Cao, Cheng Tan, Zhangyang Gao, Yilun Xu, Guangyong Chen, Pheng-Ann Heng, and Stan Z. Li. 2024. A Survey on Generative Diffusion Models. IEEE Transactions on Knowledge and Data Engineering (2024), 1–20.
- Cho et al. (2023) Jaemin Cho, Abhay Zala, and Mohit Bansal. 2023. DALL-EVAL: Probing the Reasoning Skills and Social Biases of Text-to-Image Generation Models. In ICCV. IEEE, 3020–3031.
- Choi et al. (2024) Yujin Choi, Jinseong Park, Hoki Kim, Jaewook Lee, and Saerom Park. 2024. Fair Sampling in Diffusion Models through Switching Mechanism. In AAAI. AAAI Press, 21995–22003.
- Friedrich et al. (2023) Felix Friedrich, Patrick Schramowski, Manuel Brack, Lukas Struppek, Dominik Hintersdorf, Sasha Luccioni, and Kristian Kersting. 2023. Fair Diffusion: Instructing Text-to-Image Generation Models on Fairness. CoRR abs/2302.10893 (2023).
- Goodfellow et al. (2020) Ian J. Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron C. Courville, and Yoshua Bengio. 2020. Generative adversarial networks. Commun. ACM 63, 11 (2020), 139–144.
- Harper and Konstan (2016) F. Maxwell Harper and Joseph A. Konstan. 2016. The MovieLens Datasets: History and Context. ACM Trans. Interact. Intell. Syst. 5, 4 (2016), 19:1–19:19.
- He et al. (2020) Xiangnan He, Kuan Deng, Xiang Wang, Yan Li, Yong-Dong Zhang, and Meng Wang. 2020. LightGCN: Simplifying and Powering Graph Convolution Network for Recommendation. In SIGIR. ACM, 639–648.
- He et al. (2017) Xiangnan He, Lizi Liao, Hanwang Zhang, Liqiang Nie, Xia Hu, and Tat-Seng Chua. 2017. Neural Collaborative Filtering. In WWW. ACM, 173–182.
- Ho et al. (2020) Jonathan Ho, Ajay Jain, and Pieter Abbeel. 2020. Denoising Diffusion Probabilistic Models. In NeurIPS.
- Hoogeboom et al. (2022) Emiel Hoogeboom, Victor Garcia Satorras, Clément Vignac, and Max Welling. 2022. Equivariant Diffusion for Molecule Generation in 3D. In ICML (Proceedings of Machine Learning Research, Vol. 162). PMLR, 8867–8887.
- Jiang et al. (2024) Yangqin Jiang, Yuhao Yang, Lianghao Xia, and Chao Huang. 2024. DiffKG: Knowledge Graph Diffusion Model for Recommendation. In WSDM. ACM, 313–321.
- Li et al. (2022) Xiang Li, John Thickstun, Ishaan Gulrajani, Percy Liang, and Tatsunori B. Hashimoto. 2022. Diffusion-LM Improves Controllable Text Generation. In NeurIPS.
- Li et al. (2023) Zihao Li, Aixin Sun, and Chenliang Li. 2023. DiffuRec: A Diffusion Model for Sequential Recommendation. CoRR abs/2304.00686 (2023).
- Liang et al. (2018) Dawen Liang, Rahul G. Krishnan, Matthew D. Hoffman, and Tony Jebara. 2018. Variational Autoencoders for Collaborative Filtering. In WWW. ACM, 689–698.
- Liu et al. (2023) Chengyi Liu, Wenqi Fan, Yunqing Liu, Jiatong Li, Hang Li, Hui Liu, Jiliang Tang, and Qing Li. 2023. Generative Diffusion Models on Graphs: Methods and Applications. In IJCAI. ijcai.org, 6702–6711.
- Liu et al. (2021) Qigang Liu, Lifeng Mu, Vijayan Sugumaran, Chongren Wang, and Dongmei Han. 2021. Pair-wise ranking based preference learning for points-of-interest recommendation. Knowl. Based Syst. 225 (2021), 107069.
- Ma et al. (2024) Haokai Ma, Ruobing Xie, Lei Meng, Xin Chen, Xu Zhang, Leyu Lin, and Zhanhui Kang. 2024. Plug-In Diffusion Model for Sequential Recommendation. In AAAI. AAAI Press, 8886–8894.
- Mansoury et al. (2022) Masoud Mansoury, Himan Abdollahpouri, Mykola Pechenizkiy, Bamshad Mobasher, and Robin Burke. 2022. A Graph-Based Approach for Mitigating Multi-Sided Exposure Bias in Recommender Systems. ACM Trans. Inf. Syst. 40, 2 (2022), 32:1–32:31.
- Mao et al. (2021) Kelong Mao, Jieming Zhu, Xi Xiao, Biao Lu, Zhaowei Wang, and Xiuqiang He. 2021. UltraGCN: Ultra Simplification of Graph Convolutional Networks for Recommendation. In CIKM. ACM, 1253–1262.
- Naghiaei et al. (2022) Mohammadmehdi Naghiaei, Hossein A. Rahmani, and Yashar Deldjoo. 2022. CPFair: Personalized Consumer and Producer Fairness Re-ranking for Recommender Systems. In SIGIR. ACM, 770–779.
- Niu et al. (2020) Chenhao Niu, Yang Song, Jiaming Song, Shengjia Zhao, Aditya Grover, and Stefano Ermon. 2020. Permutation Invariant Graph Generation via Score-Based Generative Modeling. In AISTATS (Proceedings of Machine Learning Research, Vol. 108). PMLR, 4474–4484.
- Qin et al. (2024) Yifang Qin, Hongjun Wu, Wei Ju, Xiao Luo, and Ming Zhang. 2024. A Diffusion Model for POI Recommendation. ACM Trans. Inf. Syst. 42, 2 (2024), 54:1–54:27.
- Rendle et al. (2009) Steffen Rendle, Christoph Freudenthaler, Zeno Gantner, and Lars Schmidt-Thieme. 2009. BPR: Bayesian Personalized Ranking from Implicit Feedback. In UAI. AUAI Press, 452–461.
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-Resolution Image Synthesis with Latent Diffusion Models. In CVPR. IEEE, 10674–10685.
- Sarwar et al. (2001) Badrul Munir Sarwar, George Karypis, Joseph A. Konstan, and John Riedl. 2001. Item-based collaborative filtering recommendation algorithms. In WWW. ACM, 285–295.
- Shen et al. (2023) Xudong Shen, Chao Du, Tianyu Pang, Min Lin, Yongkang Wong, and Mohan S. Kankanhalli. 2023. Finetuning Text-to-Image Diffusion Models for Fairness. CoRR abs/2311.07604 (2023).
- Shenbin et al. (2020) Ilya Shenbin, Anton Alekseev, Elena Tutubalina, Valentin Malykh, and Sergey I. Nikolenko. 2020. RecVAE: A New Variational Autoencoder for Top-N Recommendations with Implicit Feedback. In WSDM. ACM, 528–536.
- Steck (2019) Harald Steck. 2019. Embarrassingly Shallow Autoencoders for Sparse Data. In WWW. ACM, 3251–3257.
- Sun et al. (2020) Jianing Sun, Wei Guo, Dengcheng Zhang, Yingxue Zhang, Florence Regol, Yaochen Hu, Huifeng Guo, Ruiming Tang, Han Yuan, Xiuqiang He, and Mark Coates. 2020. A Framework for Recommending Accurate and Diverse Items Using Bayesian Graph Convolutional Neural Networks. In KDD. ACM, 2030–2039.
- Togashi et al. (2024) Riku Togashi, Kenshi Abe, and Yuta Saito. 2024. Scalable and Provably Fair Exposure Control for Large-Scale Recommender Systems. In WWW. ACM, 3307–3318.
- Vassøy and Langseth (2024) Bjørnar Vassøy and Helge Langseth. 2024. Consumer-side fairness in recommender systems: a systematic survey of methods and evaluation. Artif. Intell. Rev. 57, 4 (2024), 101.
- Wang et al. (2017) Jun Wang, Lantao Yu, Weinan Zhang, Yu Gong, Yinghui Xu, Benyou Wang, Peng Zhang, and Dell Zhang. 2017. IRGAN: A Minimax Game for Unifying Generative and Discriminative Information Retrieval Models. In SIGIR. ACM, 515–524.
- Wang et al. (2023) Wenjie Wang, Yiyan Xu, Fuli Feng, Xinyu Lin, Xiangnan He, and Tat-Seng Chua. 2023. Diffusion Recommender Model. In SIGIR. ACM, 832–841.
- Wu et al. (2023b) Likang Wu, Zhi Zheng, Zhaopeng Qiu, Hao Wang, Hongchao Gu, Tingjia Shen, Chuan Qin, Chen Zhu, Hengshu Zhu, Qi Liu, Hui Xiong, and Enhong Chen. 2023b. A Survey on Large Language Models for Recommendation. CoRR abs/2305.19860 (2023).
- Wu et al. (2023a) Shiwen Wu, Fei Sun, Wentao Zhang, Xu Xie, and Bin Cui. 2023a. Graph Neural Networks in Recommender Systems: A Survey. ACM Comput. Surv. 55, 5 (2023), 97:1–97:37.
- Xu et al. (2022) Minkai Xu, Lantao Yu, Yang Song, Chence Shi, Stefano Ermon, and Jian Tang. 2022. GeoDiff: A Geometric Diffusion Model for Molecular Conformation Generation. In ICLR. OpenReview.net.
- Yang et al. (2016) Dingqi Yang, Daqing Zhang, and Bingqing Qu. 2016. Participatory Cultural Mapping Based on Collective Behavior Data in Location-Based Social Networks. ACM Trans. Intell. Syst. Technol. 7, 3 (2016), 30:1–30:23.
- Yang et al. (2024) Ling Yang, Zhilong Zhang, Yang Song, Shenda Hong, Runsheng Xu, Yue Zhao, Wentao Zhang, Bin Cui, and Ming-Hsuan Yang. 2024. Diffusion Models: A Comprehensive Survey of Methods and Applications. ACM Comput. Surv. 56, 4 (2024), 105:1–105:39.
- Yu et al. (2024) Junliang Yu, Xin Xia, Tong Chen, Lizhen Cui, Nguyen Quoc Viet Hung, and Hongzhi Yin. 2024. XSimGCL: Towards Extremely Simple Graph Contrastive Learning for Recommendation. IEEE Trans. Knowl. Data Eng. 36, 2 (2024), 913–926.
- Zhao et al. (2022b) Minghao Zhao, Le Wu, Yile Liang, Lei Chen, Jian Zhang, Qilin Deng, Kai Wang, Xudong Shen, Tangjie Lv, and Runze Wu. 2022b. Investigating Accuracy-Novelty Performance for Graph-based Collaborative Filtering. In SIGIR. ACM, 50–59.
- Zhao et al. (2022a) Wayne Xin Zhao, Yupeng Hou, Xingyu Pan, Chen Yang, Zeyu Zhang, Zihan Lin, Jingsen Zhang, Shuqing Bian, Jiakai Tang, Wenqi Sun, Yushuo Chen, Lanling Xu, Gaowei Zhang, Zhen Tian, Changxin Tian, Shanlei Mu, Xinyan Fan, Xu Chen, and Ji-Rong Wen. 2022a. RecBole 2.0: Towards a More Up-to-Date Recommendation Library. In CIKM. ACM, 4722–4726.