### v2

# How Fair is Your Diffusion Recommender Model?

Abstract.
Diffusion-based learning has settled as a rising paradigm in generative recommendation, outperforming traditional approaches built upon variational autoencoders and generative adversarial networks. Despite their effectiveness, concerns have been raised that diffusion models - widely adopted in other machine-learning domains - could potentially lead to unfair outcomes, since they are trained to recover data distributions that often encode inherent biases. Motivated by the related literature, and acknowledging the extensive discussion around bias and fairness aspects in recommendation, we propose, to the best of our knowledge, the first empirical study of fairness for DiffRec, chronologically the pioneer technique in diffusion-based recommendation. Our empirical study involves DiffRec and its variant L-DiffRec, tested against nine recommender systems on two benchmarking datasets to assess recommendation utility and fairness from both consumer and provider perspectives. Specifically, we first evaluate the utility and fairness dimensions separately and, then, within a multi-criteria setting to investigate whether, and to what extent, these approaches can achieve a trade-off between the two. While showing worrying trends in alignment with the more general machine-learning literature on diffusion models, our results also indicate promising directions to address the unfairness issue in future work. The source code is available at https://github.com/danielemalitesta/FairDiffRec.
1. Introduction
Research in recommender systems (RSs) has recently been shaped by the latest advances in deep learning, e.g., graph neural networks (Wu et al., 2023), language models (Balloccu et al., 2023; Wu et al., 2024), and diffusion models (Wang et al., 2023). While the first two families have already received significant attention in the literature, diffusion-based RSs are still in early stages.
As of today, diffusion models (Ho et al., 2020; Yang et al., 2024) have settled as the leading solution in generative artificial intelligence (AI) (Goodfellow et al., 2020) and revolutionized the paradigm of generating new (but realistic) data with several applications, such as computer vision (Rombach et al., 2022; Ho et al., 2020), natural language processing (Austin et al., 2021; Li et al., 2022), and graph modeling (Niu et al., 2020; Liu et al., 2023), and domains, such as chemistry and bioinformatics (Xu et al., 2022; Hoogeboom et al., 2022). The core principle behind diffusion models lies in their twofold learning procedure, involving a forward process where samples from a real-data distribution are iteratively corrupted by adding noise, e.g., Gaussian noise, and a reverse process where the model is trained to reconstruct the original data by denoising the corrupted samples.
As observed by Wang et al. (2023), the diffusion model framework is well-suited for recommendation settings, as it can learn to denoise users’ implicit feedback containing noisy interactions, i.e., false positives. This approach, called DiffRec, shows improved performance with respect to traditional generative RSs, e.g., those based on variational autoencoders (VAEs) (Liang et al., 2018) and generative adversarial networks (GANs) (Wang et al., 2017), by addressing the limited representational power of the former and the unstable training procedures of the latter. Similarly, other works proposed to leverage diffusion models for specific recommendation tasks such as point-of-interests (Qin et al., 2024), sequential (Li et al., 2024b; Ma et al., 2024), or knowledge graph-based recommendation (Jiang et al., 2024).
Diffusion models have achieved widespread success in many applications and domains; however, as a relatively recent development in machine learning (ML), they continue to face several open challenges. For instance, some works have raised concerns regarding the unfair outcomes diffusion models could lead to. As observed in generative AI, bias is intrinsic in real-world data, and diffusion models might amplify this signal while learning the underlying biased data distribution. In their seminal paper, Friedrich et al. (2023) discuss how stable diffusion methods are prone to generate synthetic images that lack diversity, e.g., firefighters are only associated with white males. Other works, such as Bianchi et al. (2023), discuss how textual prompts with trait annotations may produce stereotypical images. Similarly, Cho et al. (2023) observe bias in the generated images in terms of people color tones and occupations. Interestingly, the same unfairness issues have also been observed in graph diffusion models (Vignac et al., 2023; Chen et al., 2023), where the generation of both the graph topology and the node features (Li et al., 2024a) further amplifies the natural bias already existing in the graph structure (Kose and Shen, 2024). Recent papers have proposed possible countermeasures, such as introducing fairness-guided instructions during the data generation process (Friedrich et al., 2023), obfuscating the sensitive attributes (Choi et al., 2024), or adopting an ad-hoc fine-tuning (Shen et al., 2024; Kose and Shen, 2024).
In this light, a question arises: “To what extent do diffusion models exacerbate or replicate unfairness patterns within recommender systems?” This question is pertinent given the large relevance that bias and fairness have held in recommendations for decades (Anelli et al., 2022; Mansoury et al., 2022; Sun et al., 2020; Zhao et al., 2022b; Boratto et al., 2023b). Thus, we consider it imperative to evaluate the performance of popular diffusion-based RSs rigorously, with a specific focus on potential unfairness issues, as there is currently no literature on this topic, despite a well-documented history of bias and unfairness affecting other ML-based RSs. As diffusion-based recommender models are still in their pioneer era, a thorough investigation of their performance can help identify and address observed pitfalls.
To this end, in this work, we select one of the pioneer approaches in diffusion-based recommendation, DiffRec (Wang et al., 2023) (and its variant L-DiffRec), to answer the following research question: “Is DiffRec a fair recommender system compared to the state-of-the-art?” To conduct our investigation, we consider two datasets (with users’ attribute data and binary gender label as the sensitive attribute) and nine additional state-of-the-art recommenders. Our purpose is to benchmark their performance on six recommendation metrics accounting for utility, consumer fairness (Boratto et al., 2022; Anelli et al., 2023; Vassøy and Langseth, 2024), and provider fairness (Boratto et al., 2024b; Naghiaei et al., 2022; Mansoury et al., 2022). Concretely, we run a twofold analysis, where we first monitor each dimension separately, and then in a multi-criteria setting. This twofold perspective can help identify patterns by considering all actors in the recommendation scenario (Anelli et al., 2023).
Our results depict a worrying scenario, confirming that unfairness concerns already observed for other ML tasks do indeed extend to the recommendation task, when leveraging diffusion-based approaches. Nonetheless, introducing additional components in the diffusion-based process, as for L-DiffRec, seems to mitigate the observed unfairness; indeed, these findings pave the way to novel future directions where fairness is preserved in diffusion-based recommendation. Despite its preliminary nature, to the best of our knowledge, this is the first attempt to provide a rigorous, fairness-aware evaluation of diffusion-based recommendation methods.
2. Background
This section provides the necessary notions regarding diffusion-based recommendation (with a special focus on DiffRec (Wang et al., 2023)), and consumer fairness and provider fairness in recommendation.
2.1. Diffusion Recommender Model
Building upon diffusion models, DiffRec (Wang et al., 2023) proposes to treat the recommendation task as a denoising process of historical user-item interactions, which may be overly noisy (i.e., contain false positives). In detail, the model is trained to first corrupt users’ interactions through noise (forward) and then reconstruct them via an ad-hoc denoising procedure (reverse). Unlike traditional diffusion models for images, users’ historical interactions are not directly corrupted into pure noise, but the ratio of added noise decreases consistently.
Let and be the sets of users and items in the system, respectively, with as the vector of user-item interactions for user . Each vector entry denotes user interacted with item , otherwise . Then, the forward process, in a Markov chain process, corrupts the user interaction vector into steps by gradually adding Gaussian noise. The formulation for any step is: , where (with a notation abuse) we indicate , , at the initial step, stands for the Gaussian distribution, controls the magnitude of the Gaussian noise added at each step , and is the standard deviation of the distribution. Conversely, the reverse process involves training any ML model, e.g., a neural network, to approximate the mean and standard deviation of the data distribution and recover the original user interaction vector. Again, for any step , we have: , where and are the learned mean and standard deviation through the ML model parametrized by .
Furthermore, to address large-scale recommendation (a pretty common scenario in real-world applications), Wang et al. (2023) also propose L-DiffRec, a lighter-weight version of DiffRec. This variant first clusters pre-trained item embeddings, e.g., from LightGCN, into clusters. Afterward, the initial user interaction vector is split into sub-vectors according to the corresponding -th item cluster. Then, a VAE is trained to compress each user interaction sub-vector , so that the diffusion, i.e., forward/reverse processes, occurs in a low-dimensional latent space, ideally in parallel computation across all splits. Finally, the original user interaction vector is reconstructed through the decoding process of the previous VAE.
Note that in Wang et al. (2023), the authors eventually propose a second variant to DiffRec, called T-DiffRec, which can encode the temporal information during the training. However, T-DiffRec is specialized for sequential recommendation, which would require a dedicated study to compare its performance with other analogous models. To this end, we exclude T-DiffRec from our experiments, leaving its for future extensions of our study.
Our analysis aims to offer a preliminary outline on the fairness performance of diffusion-based RSs. Although other works have recently contributed with novel diffusion-based solutions (Qin et al., 2024; Li et al., 2024b; Ma et al., 2024; Jiang et al., 2024), DiffRec could be rightly considered as the pioneer approach in this domain for chronological reasons. In light of this, we plan to extend our analysis to other diffusion models in future works.
2.2. Fairness Evaluation in Recommendation
Recent studies, both in RS (Boratto et al., 2022; Naghiaei et al., 2022; Vassøy and Langseth, 2024; Boratto et al., 2023a; Medda et al., 2025) and relatedly user modeling (Purificato et al., 2025, 2022, 2024), focused on assessing algorithmic fairness under the notion of demographic parity (Dwork et al., 2012; Feldman et al., 2015). From a consumer perspective (Boratto et al., 2024a; Naghiaei et al., 2022), we operationalize this notion as the absolute difference in recommendation utility between users’ groups. To this end, we select the respective fairness counterparts of nDCG and Recall, namely nDCG and Recall. On the other hand, provider fairness is estimated as the absolute difference in exposure (i.e., how evenly items or groups of items are exposed to users or groups of users) between items’ groups (Exp) (Naghiaei et al., 2022; Boratto et al., 2024b) and as the average percentage of long-tail items in the top- lists (APLT) (Anelli et al., 2023; Mansoury et al., 2022). For all metrics except APLT, values closer to zero indicate fairer recommendations. Due to space constraints, we do not delve into the detailed formulations. Interested readers are referred to the cited works for a comprehensive presentation.
3. Experimental Study
In this section, we present our experimental study aimed at assessing recommendation utility and fairness in diffusion-based methods. First, we describe the baselines and datasets adopted for the analysis. Then, we provide reproducibility details for the experiments.
3.1. Baselines
We select nine state-of-the-art recommendation models along with DiffRec and L-DiffRec: BPRMF (Rendle et al., 2009), ItemkNN (Sarwar et al., 2001), NeuMF (He et al., 2017), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), XSimGCL (Yu et al., 2024), EASE (Steck, 2019), MultiVAE (Liang et al., 2018), and RecVAE (Shenbin et al., 2020). As reported in Table 1, our selection was guided by the following observations: (i) the baselines are well-established solutions, accepted at top-tier conferences and journals; (ii) they cover a diverse set of recommendation families, such as graph neural networks and generative models; (iii) many of these approaches have been already analyzed from a recommendation fairness perspective.
3.2. Datasets
We assess the performance of the selected baselines on MovieLens-1M (ML1M) (Harper and Konstan, 2016) and Foursquare Tokyo (FTKY) (Yang et al., 2016), which are frequently adopted in fairness-aware literature in recommendation (Vassøy and Langseth, 2024; Boratto et al., 2024a; Liu et al., 2021). While ML1M is a popular recommendation dataset that collects movies and users’ preferences on them, FTKY includes users’ check-ins at points of interest in Tokyo. In both datasets, users’ sensitive information comes in a binary gender label. Thus, fairness analyses on the consumer side adopt the gender attribute included in both datasets (where female is the protected group), whereas provider fairness is assessed on a binary partition of the items’ set, namely short-head (popular) and long-tail (niche) items (where niche items are the protected group). Following a standard procedure in the literature, as in (Boratto et al., 2024b; Naghiaei et al., 2022; Mansoury et al., 2022), we select the 20% most popular items as short-head and the rest as long-tail. Dataset statistics are reported in Table 2.
3.3. Reproducibility Details
Datasets are split into train, validation, and test sets (70%/10%/20%) following the temporal hold-out method adopted in (Wang et al., 2023; Boratto et al., 2024a). We use the RecBole framework (Zhao et al., 2022a) to train and evaluate all baseline models. For the two diffusion-based recommender systems, we adopt their original code implementations111https://github.com/YiyanXu/DiffRec. Accessed August 10, 2025.. Baseline models are trained using a grid search over hyperparameters, guided by the settings in the original papers. The configuration yielding the highest Recall@20 on the validation set is selected for each model.
| Models | Year | Venue | Fairness analyzed in |
| ItemkNN (Sarwar et al., 2001) | 2001 | WWW | (Anelli et al., 2022; Boratto et al., 2023b, 2022) |
| BPRMF (Rendle et al., 2009) | 2009 | UAI | (Anelli et al., 2022, 2023; Mansoury et al., 2022; Sun et al., 2020; Zhao et al., 2022b; Boratto et al., 2022) |
| NeuMF (He et al., 2017) | 2017 | WWW | (Anelli et al., 2022; Boratto et al., 2023b; Mansoury et al., 2022; Sun et al., 2020; Boratto et al., 2022) |
| MultiVAE (Liang et al., 2018) | 2018 | WWW | (Anelli et al., 2022; Togashi et al., 2024) |
| EASE (Steck, 2019) | 2019 | WWW | (Anelli et al., 2022) |
| RecVAE (Shenbin et al., 2020) | 2020 | WSDM | ✗ |
| LightGCN (He et al., 2020) | 2020 | SIGIR | (Anelli et al., 2023; Boratto et al., 2024a; Zhao et al., 2022b) |
| UltraGCN (Mao et al., 2021) | 2021 | CIKM | (Anelli et al., 2023; Boratto et al., 2024a) |
| DiffRec (Wang et al., 2023) | 2023 | SIGIR | ✗ |
| L-DiffRec (Wang et al., 2023) | 2023 | SIGIR | ✗ |
| XSimGCL (Yu et al., 2024) | 2024 | TKDE | (Boratto et al., 2024a) |
| Datasets | ML1M (Harper and Konstan, 2016) | FTKY (Yang et al., 2016) |
| # Users | 6,040 | 7,240 |
| # Items | 3,650 | 5,779 |
| # Interactions | 1,000,130 | 353,769 |
| Sparsity | 95.46% | 99.15% |
| Gender Repr. | M : 71.7%; F : 28.3% | M : 87.9%; F : 12.1% |
| Min. Interactions | 20 | 20 |
4. Results and Discussion
In this section, we aim to answer the following research question: “Is DiffRec a fair RS compared to state-of-the-art methods?”. To this end, we first investigate the recommendation utility and fairness dimensions separately, and then we inspect the trade-off among metrics from both dimensions to assess whether DiffRec can deliver recommendations that are both fair and effective.
| Metrics | Random | Pop | BPRMF | ItemkNN | NeuMF | LightGCN | UltraGCN | XSimGCL | EASE | MultiVAE | RecVAE | DiffRec | L-DiffRec |
| Movielens-1M (ML1M) | |||||||||||||
| Recall () | 0.61 | 6.88 | 10.67 | 9.87 | 10.44 | 10.16 | 10.74 | 10.46 | 6.45 | 8.25 | 9.68 | 10.71 | 10.38 |
| nDCG () | 1.14 | 10.42 | 12.94 | 12.22 | 12.67 | 13.09 | 12.82 | 12.80 | 8.62 | 12.04 | 12.86 | 13.19 | 12.85 |
| Recall () | 0.01 | 0.82 | 0.43 | 1.08 | 0.32 | 0.19 | 0.06 | 0.14 | 0.70 | 0.48 | 0.53 | 0.86 | 0.66 |
| nDCG () | 0.19 | 2.54 | 1.57 | 1.98 | 1.54 | 1.43 | 1.37 | 1.34 | 1.66 | 2.43 | 2.07 | 2.03 | 2.25 |
| APLT () | 79.80 | 0.00 | 12.47 | 8.01 | 11.75 | 8.11 | 13.73 | 10.50 | 9.00 | 2.63 | 5.38 | 6.18 | 10.81 |
| Exp () | 0.12 | 100.00 | 86.29 | 91.18 | 87.36 | 91.32 | 84.98 | 88.78 | 90.54 | 97.02 | 94.10 | 93.67 | 87.26 |
| Fousquare Tokyo (FTKY) | |||||||||||||
| Recall () | 0.30 | 7.89 | 12.17 | 11.77 | 12.48 | 11.03 | 12.22 | 13.07 | 6.19 | 10.06 | 12.30 | 13.10 | 12.48 |
| nDCG () | 0.22 | 5.57 | 10.73 | 10.48 | 10.93 | 9.87 | 10.85 | 11.37 | 5.48 | 9.21 | 10.90 | 11.53 | 10.77 |
| Recall () | 0.03 | 0.95 | 0.94 | 0.69 | 0.59 | 0.77 | 1.05 | 0.98 | 0.40 | 0.97 | 1.00 | 0.80 | 0.61 |
| nDCG () | 0.01 | 0.85 | 0.22 | 0.19 | 0.09 | 0.47 | 0.29 | 0.23 | 0.10 | 0.59 | 0.29 | 0.22 | 0.02 |
| APLT () | 80.00 | 0.00 | 9.93 | 18.27 | 10.20 | 7.42 | 5.36 | 11.06 | 6.32 | 3.72 | 7.27 | 10.75 | 16.97 |
| Exp () | 0.09 | 100.00 | 89.27 | 78.91 | 89.39 | 91.87 | 94.32 | 88.25 | 93.56 | 95.75 | 92.03 | 88.57 | 80.85 |
4.1. Uni-Dimensional Analysis
Table 3 shows the results of the selected baselines (along with reference models such as Random and Pop) on ML1M and FTKY in terms of recommendation utility (i.e., Recall and nDCG), consumer fairness (i.e., Recall and nDCG), and provider fairness (i.e., APLT and Exp). Note that each metric is computed on the top-20 recommendation lists and is expressed as a percentage (%).
On the recommendation utility dimension, the results are closely aligned with those observed in the related literature, where graph- and diffusion-based RSs generally dominate over the other baselines on both recommendation datasets. DiffRec shows slightly superior performance than its variant L-DiffRec. However, it should be noted that the latter is more computationally lightweight. This gap between the two DiffRec variants might be due to the diffusion process being conducted in the latent space by L-DiffRec. The VAEs employed by L-DiffRec compress the users’ preferences signals, potentially leading to information loss and non-perfect reconstruction during decoding. Conversely, the DiffRec forward and reverse processes fully capture the discrete preference information, without dealing with the potential loss caused by information bottlenecks.
Moving to the experiments on the fairness dimension, our results exhibit distinct trends across the datasets. On ML1M, graph-based approaches such as UltraGCN and XSimGCL can provide both highly effective and fair recommendations, while the diffusion-based models settle among the least fair approaches among the considered methods. On FTKY, the fairest RSs are instead the generative ones, such as EASE and L-DiffRec, and the traditional ones, such as ItemKNN and NeuMF. The diverse trends in consumer fairness across ML1M and FTKY might derive from their different domains. Specifically, users’ interactions in the movie domain are driven by intrinsically individual preferences towards specific movie categories, such as those sharing the same actors or the same genre. In contrast, venue check-ins in FTKY might derive from going out decided as groups, e.g., with friends or partners. Biases in movie preferences may be more strongly amplified by recommender systems compared to biases in venue preferences.
The consumer unfairness exhibited by DiffRec on ML1M shows that diffusion models are prone to amplify biases, as also reported by other studies. However, a common observation across the datasets is that the L-DiffRec variant leads to fairer recommendations than DiffRec and the other considered generative models, e.g., MultiVAE and RecVAE. Particularly, regarding APLT and Exp, L-DiffRec exhibits fairer outcomes on the provider side. This might be attributed to the intrinsic clustering operation of L-DiffRec, which shifts the learning process from following the common popularity bias to capturing more nuanced patterns across items’ categories.
4.2. Multi-Dimensional Analysis
As the three analyzed properties, i.e., recommendation utility, consumer fairness, and provider fairness, represent different and possibly contrasting dimensions of the recommendation scenario, we extend our previous analysis by assessing the trade-off among them, as done in related studies (Anelli et al., 2023; Boratto et al., 2023b). Indeed, we believe that this additional assessment may complement the findings from the uni-dimensional analysis, motivating and explaining from other complementary perspectives the previous performance trends.
The Kiviat diagrams in Figure 1 represent the performance for both datasets of five representative baselines, i.e., ItemkNN, XSimGCL, RecVAE, DiffRec, and L-DiffRec, on three metrics, i.e., nDCG, nDCG, and APLT, accounting for utility, consumer, and provider fairness. To provide an easier interpretation of the plots, we follow (Anelli et al., 2023) and normalize the metrics into [0, 1] through min-max normalization and rescale nDCG to adhere to the principle that “the higher the values the better the performance”.
On the ML1M dataset, the findings from the uni-dimensional analysis are generally confirmed. Indeed, we observe that XSimGCL is the most balanced baseline across all models, being capable of providing an optimal trade-off among all metrics. Conversely, ItemkNN, RecVAE, and DiffRec cannot generally increase one metric side without harming the other ones. Interestingly, though, while ItemkNN tends to maximize fairness over the utility, RecVAE and DiffRec, i.e., generative and diffusion-based models, do the opposite. This becomes especially evident on DiffRec, where the model is largely improving utility at the detriment of provider/consumer fairness. The exception to the trend is L-DiffRec, reaching the second-best trade-off soon after XSimGCL. In fact, on FTKY, L-DiffRec exhibits even the best trade-off compared to the selected baselines.
In summary, the findings from the multi-dimensional analysis confirm the unfairness warnings raised in the uni-dimensional analysis but, at the same time, suggest that diffusion-based methods may have room for improvement in achieving fairer recommendations.
5. Conclusion and Future Work
This work serves as a preliminary attempt to rigorously assess the performance of diffusion-based RSs, specifically focusing on one of the pioneer approaches, DiffRec, and its variant L-DiffRec. The relevance of our study is underscored by the ML literature, which has recently noted that the intrinsic bias within the data may lead to diffusion models generating unfair outcomes. To this end, we decided to tailor our investigation to the performance of (L-)DiffRec under common fairness properties in RSs. Specifically, we conducted an extensive analysis involving two recommendation datasets (with sensitive information), nine state-of-the-art RSs, and six recommendation metrics that account for accuracy, consumer, and provider fairness.
Results in two complementary evaluation settings, conducted on single metrics and in a trade-off scenario, empirically confirmed that the same bias amplification caused by diffusion models in ML seems to be advisable also in the recommendation domain (DiffRec). However, when adopting diffusion strategies carefully designed to tailor the recommendation scenario (L-DiffRec), the unfairness issue can be solved and even reverted. Thus, our empirical study suggests that diffusion models cannot be plugged into RSs as they are, since they would bring (and even exacerbate) the unfairness issues already observed in ML.
However, with careful and tailored modifications to the rationale and models’ architecture (e.g., operating in a clustered latent space, as done in L-DiffRec, to reduce noise sensitivity and group-level bias), the issues might be properly addressed.
To confirm our final assumption, future works will aim to extend our study to the other DiffRec variant, T-DiffRec, as well as other diffusion-based RSs (Walker et al., 2022; Hou et al., 2024; Choi et al., 2024). While trying to confirm the same observed (un)fairness trends as in this work, we will also devote future investigations to understanding the source of unfairness more carefully, for instance, by examining how the recommendation dataset is modified after the generation process of the diffusion model. Finally, on such theoretical and empirical validations, we plan to exploit the lessons learned from this empirical study and propose tailored strategies to mitigate (when present) the unfairness of diffusion-based recommendation systems.
Acknowledgements.
Daniele Malitesta and Fragkiskos D. Malliaros acknowledge the support of the Innov4-ePiK project managed by the French National Research Agency under the 4th PIA, integrated into France2030 (ANR-23-RHUS-0002).References
- (1)
- Anelli et al. (2022) Vito Walter Anelli, Alejandro Bellogín, Tommaso Di Noia, Dietmar Jannach, and Claudio Pomo. 2022. Top-N Recommendation Algorithms: A Quest for the State-of-the-Art. In UMAP ’22: 30th ACM Conference on User Modeling, Adaptation and Personalization, Barcelona, Spain, July 4 - 7, 2022, Alejandro Bellogín, Ludovico Boratto, Olga C. Santos, Liliana Ardissono, and Bart P. Knijnenburg (Eds.). ACM, 121–131. https://doi.org/10.1145/3503252.3531292
- Anelli et al. (2023) Vito Walter Anelli, Yashar Deldjoo, Tommaso Di Noia, Daniele Malitesta, Vincenzo Paparella, and Claudio Pomo. 2023. Auditing Consumer- and Producer-Fairness in Graph Collaborative Filtering. In Advances in Information Retrieval - 45th European Conference on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2-6, 2023, Proceedings, Part I (Lecture Notes in Computer Science, Vol. 13980), Jaap Kamps, Lorraine Goeuriot, Fabio Crestani, Maria Maistro, Hideo Joho, Brian Davis, Cathal Gurrin, Udo Kruschwitz, and Annalina Caputo (Eds.). Springer, 33–48. https://doi.org/10.1007/978-3-031-28244-7_3
- Austin et al. (2021) Jacob Austin, Daniel D. Johnson, Jonathan Ho, Daniel Tarlow, and Rianne van den Berg. 2021. Structured Denoising Diffusion Models in Discrete State-Spaces. In Advances in Neural Information Processing Systems 34: Annual Conference on Neural Information Processing Systems 2021, NeurIPS 2021, December 6-14, 2021, virtual, Marc’Aurelio Ranzato, Alina Beygelzimer, Yann N. Dauphin, Percy Liang, and Jennifer Wortman Vaughan (Eds.). 17981–17993. https://proceedings.neurips.cc/paper/2021/hash/958c530554f78bcd8e97125b70e6973d-Abstract.html
- Balloccu et al. (2023) Giacomo Balloccu, Ludovico Boratto, Christian Cancedda, Gianni Fenu, and Mirko Marras. 2023. Knowledge is Power, Understanding is Impact: Utility and Beyond Goals, Explanation Quality, and Fairness in Path Reasoning Recommendation. In Advances in Information Retrieval - 45th European Conference on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2-6, 2023, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 13982), Jaap Kamps, Lorraine Goeuriot, Fabio Crestani, Maria Maistro, Hideo Joho, Brian Davis, Cathal Gurrin, Udo Kruschwitz, and Annalina Caputo (Eds.). Springer, 3–19. https://doi.org/10.1007/978-3-031-28241-6_1
- Bianchi et al. (2023) Federico Bianchi, Pratyusha Kalluri, Esin Durmus, Faisal Ladhak, Myra Cheng, Debora Nozza, Tatsunori Hashimoto, Dan Jurafsky, James Zou, and Aylin Caliskan. 2023. Easily Accessible Text-to-Image Generation Amplifies Demographic Stereotypes at Large Scale. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency, FAccT 2023, Chicago, IL, USA, June 12-15, 2023. ACM, 1493–1504. https://doi.org/10.1145/3593013.3594095
- Boratto et al. (2023a) Ludovico Boratto, Francesco Fabbri, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2023a. Counterfactual Graph Augmentation for Consumer Unfairness Mitigation in Recommender Systems. In Proceedings of the 32nd ACM International Conference on Information and Knowledge Management, CIKM 2023, Birmingham, United Kingdom, October 21-25, 2023, Ingo Frommholz, Frank Hopfgartner, Mark Lee, Michael Oakes, Mounia Lalmas, Min Zhang, and Rodrygo L. T. Santos (Eds.). ACM, 3753–3757. https://doi.org/10.1145/3583780.3615165
- Boratto et al. (2024a) Ludovico Boratto, Francesco Fabbri, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2024a. Fair Augmentation for Graph Collaborative Filtering. In Proceedings of the 18th ACM Conference on Recommender Systems, RecSys 2024, Bari, Italy, October 14-18, 2024, Tommaso Di Noia, Pasquale Lops, Thorsten Joachims, Katrien Verbert, Pablo Castells, Zhenhua Dong, and Ben London (Eds.). ACM, 158–168. https://doi.org/10.1145/3640457.3688064
- Boratto et al. (2024b) Ludovico Boratto, Francesco Fabbri, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2024b. Robustness in Fairness Against Edge-Level Perturbations in GNN-Based Recommendation. In Advances in Information Retrieval - 46th European Conference on Information Retrieval, ECIR 2024, Glasgow, UK, March 24-28, 2024, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 14610), Nazli Goharian, Nicola Tonellotto, Yulan He, Aldo Lipani, Graham McDonald, Craig Macdonald, and Iadh Ounis (Eds.). Springer, 38–55. https://doi.org/10.1007/978-3-031-56063-7_3
- Boratto et al. (2022) Ludovico Boratto, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2022. Consumer Fairness in Recommender Systems: Contextualizing Definitions and Mitigations. In Advances in Information Retrieval - 44th European Conference on IR Research, ECIR 2022, Stavanger, Norway, April 10-14, 2022, Proceedings, Part I (Lecture Notes in Computer Science, Vol. 13185), Matthias Hagen, Suzan Verberne, Craig Macdonald, Christin Seifert, Krisztian Balog, Kjetil Nørvåg, and Vinay Setty (Eds.). Springer, 552–566. https://doi.org/10.1007/978-3-030-99736-6_37
- Boratto et al. (2023b) Ludovico Boratto, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2023b. Practical perspectives of consumer fairness in recommendation. Inf. Process. Manag. 60, 2 (2023), 103208. https://doi.org/10.1016/J.IPM.2022.103208
- Chen et al. (2023) Xiaohui Chen, Jiaxing He, Xu Han, and Liping Liu. 2023. Efficient and Degree-Guided Graph Generation via Discrete Diffusion Modeling. In International Conference on Machine Learning, ICML 2023, 23-29 July 2023, Honolulu, Hawaii, USA (Proceedings of Machine Learning Research, Vol. 202), Andreas Krause, Emma Brunskill, Kyunghyun Cho, Barbara Engelhardt, Sivan Sabato, and Jonathan Scarlett (Eds.). PMLR, 4585–4610. https://proceedings.mlr.press/v202/chen23k.html
- Cho et al. (2023) Jaemin Cho, Abhay Zala, and Mohit Bansal. 2023. DALL-EVAL: Probing the Reasoning Skills and Social Biases of Text-to-Image Generation Models. In IEEE/CVF International Conference on Computer Vision, ICCV 2023, Paris, France, October 1-6, 2023. IEEE, 3020–3031. https://doi.org/10.1109/ICCV51070.2023.00283
- Choi et al. (2024) Yujin Choi, Jinseong Park, Hoki Kim, Jaewook Lee, and Saerom Park. 2024. Fair Sampling in Diffusion Models through Switching Mechanism. In Thirty-Eighth AAAI Conference on Artificial Intelligence, AAAI 2024, Thirty-Sixth Conference on Innovative Applications of Artificial Intelligence, IAAI 2024, Fourteenth Symposium on Educational Advances in Artificial Intelligence, EAAI 2014, February 20-27, 2024, Vancouver, Canada, Michael J. Wooldridge, Jennifer G. Dy, and Sriraam Natarajan (Eds.). AAAI Press, 21995–22003. https://doi.org/10.1609/AAAI.V38I20.30202
- Dwork et al. (2012) Cynthia Dwork, Moritz Hardt, Toniann Pitassi, Omer Reingold, and Richard S. Zemel. 2012. Fairness through awareness. In Innovations in Theoretical Computer Science 2012, Cambridge, MA, USA, January 8-10, 2012, Shafi Goldwasser (Ed.). ACM, 214–226. https://doi.org/10.1145/2090236.2090255
- Feldman et al. (2015) Michael Feldman, Sorelle A. Friedler, John Moeller, Carlos Scheidegger, and Suresh Venkatasubramanian. 2015. Certifying and Removing Disparate Impact. In Proceedings of the 21th ACM SIGKDD International Conference on Knowledge Discovery and Data Mining, Sydney, NSW, Australia, August 10-13, 2015, Longbing Cao, Chengqi Zhang, Thorsten Joachims, Geoffrey I. Webb, Dragos D. Margineantu, and Graham Williams (Eds.). ACM, 259–268. https://doi.org/10.1145/2783258.2783311
- Friedrich et al. (2023) Felix Friedrich, Patrick Schramowski, Manuel Brack, Lukas Struppek, Dominik Hintersdorf, Sasha Luccioni, and Kristian Kersting. 2023. Fair Diffusion: Instructing Text-to-Image Generation Models on Fairness. CoRR abs/2302.10893 (2023). https://doi.org/10.48550/ARXIV.2302.10893 arXiv:2302.10893
- Goodfellow et al. (2020) Ian J. Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron C. Courville, and Yoshua Bengio. 2020. Generative adversarial networks. Commun. ACM 63, 11 (2020), 139–144. https://doi.org/10.1145/3422622
- Harper and Konstan (2016) F. Maxwell Harper and Joseph A. Konstan. 2016. The MovieLens Datasets: History and Context. ACM Trans. Interact. Intell. Syst. 5, 4 (2016), 19:1–19:19. https://doi.org/10.1145/2827872
- He et al. (2020) Xiangnan He, Kuan Deng, Xiang Wang, Yan Li, Yong-Dong Zhang, and Meng Wang. 2020. LightGCN: Simplifying and Powering Graph Convolution Network for Recommendation. In Proceedings of the 43rd International ACM SIGIR conference on research and development in Information Retrieval, SIGIR 2020, Virtual Event, China, July 25-30, 2020, Jimmy X. Huang, Yi Chang, Xueqi Cheng, Jaap Kamps, Vanessa Murdock, Ji-Rong Wen, and Yiqun Liu (Eds.). ACM, 639–648. https://doi.org/10.1145/3397271.3401063
- He et al. (2017) Xiangnan He, Lizi Liao, Hanwang Zhang, Liqiang Nie, Xia Hu, and Tat-Seng Chua. 2017. Neural Collaborative Filtering. In Proceedings of the 26th International Conference on World Wide Web, WWW 2017, Perth, Australia, April 3-7, 2017, Rick Barrett, Rick Cummings, Eugene Agichtein, and Evgeniy Gabrilovich (Eds.). ACM, 173–182. https://doi.org/10.1145/3038912.3052569
- Ho et al. (2020) Jonathan Ho, Ajay Jain, and Pieter Abbeel. 2020. Denoising Diffusion Probabilistic Models. In Advances in Neural Information Processing Systems 33: Annual Conference on Neural Information Processing Systems 2020, NeurIPS 2020, December 6-12, 2020, virtual, Hugo Larochelle, Marc’Aurelio Ranzato, Raia Hadsell, Maria-Florina Balcan, and Hsuan-Tien Lin (Eds.). https://proceedings.neurips.cc/paper/2020/hash/4c5bcfec8584af0d967f1ab10179ca4b-Abstract.html
- Hoogeboom et al. (2022) Emiel Hoogeboom, Victor Garcia Satorras, Clément Vignac, and Max Welling. 2022. Equivariant Diffusion for Molecule Generation in 3D. In International Conference on Machine Learning, ICML 2022, 17-23 July 2022, Baltimore, Maryland, USA (Proceedings of Machine Learning Research, Vol. 162), Kamalika Chaudhuri, Stefanie Jegelka, Le Song, Csaba Szepesvári, Gang Niu, and Sivan Sabato (Eds.). PMLR, 8867–8887. https://proceedings.mlr.press/v162/hoogeboom22a.html
- Hou et al. (2024) Yu Hou, Jin-Duk Park, and Won-Yong Shin. 2024. Collaborative Filtering Based on Diffusion Models: Unveiling the Potential of High-Order Connectivity. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2024, Washington DC, USA, July 14-18, 2024, Grace Hui Yang, Hongning Wang, Sam Han, Claudia Hauff, Guido Zuccon, and Yi Zhang (Eds.). ACM, 1360–1369. https://doi.org/10.1145/3626772.3657742
- Jiang et al. (2024) Yangqin Jiang, Yuhao Yang, Lianghao Xia, and Chao Huang. 2024. DiffKG: Knowledge Graph Diffusion Model for Recommendation. In Proceedings of the 17th ACM International Conference on Web Search and Data Mining, WSDM 2024, Merida, Mexico, March 4-8, 2024, Luz Angelica Caudillo-Mata, Silvio Lattanzi, Andrés Muñoz Medina, Leman Akoglu, Aristides Gionis, and Sergei Vassilvitskii (Eds.). ACM, 313–321. https://doi.org/10.1145/3616855.3635850
- Kose and Shen (2024) Oyku Deniz Kose and Yanning Shen. 2024. FairWire: Fair Graph Generation. In Advances in Neural Information Processing Systems 38: Annual Conference on Neural Information Processing Systems 2024, NeurIPS 2024, Vancouver, BC, Canada, December 10 - 15, 2024, Amir Globersons, Lester Mackey, Danielle Belgrave, Angela Fan, Ulrich Paquet, Jakub M. Tomczak, and Cheng Zhang (Eds.). http://papers.nips.cc/paper_files/paper/2024/hash/e105d1cabb6ee2495595e2baf25493b5-Abstract-Conference.html
- Li et al. (2024a) Mufei Li, Eleonora Kreacic, Vamsi K. Potluru, and Pan Li. 2024a. GraphMaker: Can Diffusion Models Generate Large Attributed Graphs? Trans. Mach. Learn. Res. 2024 (2024). https://openreview.net/forum?id=0q4zjGMKoA
- Li et al. (2022) Xiang Lisa Li, John Thickstun, Ishaan Gulrajani, Percy Liang, and Tatsunori B. Hashimoto. 2022. Diffusion-LM Improves Controllable Text Generation. In Advances in Neural Information Processing Systems 35: Annual Conference on Neural Information Processing Systems 2022, NeurIPS 2022, New Orleans, LA, USA, November 28 - December 9, 2022, Sanmi Koyejo, S. Mohamed, A. Agarwal, Danielle Belgrave, K. Cho, and A. Oh (Eds.). http://papers.nips.cc/paper_files/paper/2022/hash/1be5bc25d50895ee656b8c2d9eb89d6a-Abstract-Conference.html
- Li et al. (2024b) Zihao Li, Aixin Sun, and Chenliang Li. 2024b. DiffuRec: A Diffusion Model for Sequential Recommendation. ACM Trans. Inf. Syst. 42, 3 (2024), 66:1–66:28. https://doi.org/10.1145/3631116
- Liang et al. (2018) Dawen Liang, Rahul G. Krishnan, Matthew D. Hoffman, and Tony Jebara. 2018. Variational Autoencoders for Collaborative Filtering. In Proceedings of the 2018 World Wide Web Conference on World Wide Web, WWW 2018, Lyon, France, April 23-27, 2018, Pierre-Antoine Champin, Fabien Gandon, Mounia Lalmas, and Panagiotis G. Ipeirotis (Eds.). ACM, 689–698. https://doi.org/10.1145/3178876.3186150
- Liu et al. (2023) Chengyi Liu, Wenqi Fan, Yunqing Liu, Jiatong Li, Hang Li, Hui Liu, Jiliang Tang, and Qing Li. 2023. Generative Diffusion Models on Graphs: Methods and Applications. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence, IJCAI 2023, 19th-25th August 2023, Macao, SAR, China. ijcai.org, 6702–6711. https://doi.org/10.24963/IJCAI.2023/751
- Liu et al. (2021) Qigang Liu, Lifeng Mu, Vijayan Sugumaran, Chongren Wang, and Dongmei Han. 2021. Pair-wise ranking based preference learning for points-of-interest recommendation. Knowl. Based Syst. 225 (2021), 107069. https://doi.org/10.1016/J.KNOSYS.2021.107069
- Ma et al. (2024) Haokai Ma, Ruobing Xie, Lei Meng, Xin Chen, Xu Zhang, Leyu Lin, and Zhanhui Kang. 2024. Plug-In Diffusion Model for Sequential Recommendation. In Thirty-Eighth AAAI Conference on Artificial Intelligence, AAAI 2024, Thirty-Sixth Conference on Innovative Applications of Artificial Intelligence, IAAI 2024, Fourteenth Symposium on Educational Advances in Artificial Intelligence, EAAI 2014, February 20-27, 2024, Vancouver, Canada, Michael J. Wooldridge, Jennifer G. Dy, and Sriraam Natarajan (Eds.). AAAI Press, 8886–8894. https://doi.org/10.1609/AAAI.V38I8.28736
- Mansoury et al. (2022) Masoud Mansoury, Himan Abdollahpouri, Mykola Pechenizkiy, Bamshad Mobasher, and Robin Burke. 2022. A Graph-Based Approach for Mitigating Multi-Sided Exposure Bias in Recommender Systems. ACM Trans. Inf. Syst. 40, 2 (2022), 32:1–32:31. https://doi.org/10.1145/3470948
- Mao et al. (2021) Kelong Mao, Jieming Zhu, Xi Xiao, Biao Lu, Zhaowei Wang, and Xiuqiang He. 2021. UltraGCN: Ultra Simplification of Graph Convolutional Networks for Recommendation. In CIKM ’21: The 30th ACM International Conference on Information and Knowledge Management, Virtual Event, Queensland, Australia, November 1 - 5, 2021, Gianluca Demartini, Guido Zuccon, J. Shane Culpepper, Zi Huang, and Hanghang Tong (Eds.). ACM, 1253–1262. https://doi.org/10.1145/3459637.3482291
- Medda et al. (2025) Giacomo Medda, Francesco Fabbri, Mirko Marras, Ludovico Boratto, and Gianni Fenu. 2025. GNNUERS: Fairness Explanation in GNNs for Recommendation via Counterfactual Reasoning. ACM Trans. Intell. Syst. Technol. 16, 1 (2025), 6:1–6:26. https://doi.org/10.1145/3655631
- Naghiaei et al. (2022) Mohammadmehdi Naghiaei, Hossein A. Rahmani, and Yashar Deldjoo. 2022. CPFair: Personalized Consumer and Producer Fairness Re-ranking for Recommender Systems. In SIGIR ’22: The 45th International ACM SIGIR Conference on Research and Development in Information Retrieval, Madrid, Spain, July 11 - 15, 2022, Enrique Amigó, Pablo Castells, Julio Gonzalo, Ben Carterette, J. Shane Culpepper, and Gabriella Kazai (Eds.). ACM, 770–779. https://doi.org/10.1145/3477495.3531959
- Niu et al. (2020) Chenhao Niu, Yang Song, Jiaming Song, Shengjia Zhao, Aditya Grover, and Stefano Ermon. 2020. Permutation Invariant Graph Generation via Score-Based Generative Modeling. In The 23rd International Conference on Artificial Intelligence and Statistics, AISTATS 2020, 26-28 August 2020, Online [Palermo, Sicily, Italy] (Proceedings of Machine Learning Research, Vol. 108), Silvia Chiappa and Roberto Calandra (Eds.). PMLR, 4474–4484. http://proceedings.mlr.press/v108/niu20a.html
- Purificato et al. (2022) Erasmo Purificato, Ludovico Boratto, and Ernesto William De Luca. 2022. Do Graph Neural Networks Build Fair User Models? Assessing Disparate Impact and Mistreatment in Behavioural User Profiling. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management, Atlanta, GA, USA, October 17-21, 2022, Mohammad Al Hasan and Li Xiong (Eds.). ACM, 4399–4403. https://doi.org/10.1145/3511808.3557584
- Purificato et al. (2024) Erasmo Purificato, Ludovico Boratto, and Ernesto William De Luca. 2024. Toward a Responsible Fairness Analysis: From Binary to Multiclass and Multigroup Assessment in Graph Neural Network-Based User Modeling Tasks. Minds Mach. 34, 3 (2024), 33. https://doi.org/10.1007/S11023-024-09685-X
- Purificato et al. (2025) Erasmo Purificato, Hannan Javed Mahadik, Ludovico Boratto, and Ernesto William De Luca. 2025. GNN’s FAME: Fairness-Aware MEssages for Graph Neural Networks. In Proceedings of the 33rd ACM Conference on User Modeling, Adaptation and Personalization, UMAP 2025, New York City, NY, USA, June 16-19, 2025. ACM, 301–306. https://doi.org/10.1145/3699682.3728324
- Qin et al. (2024) Yifang Qin, Hongjun Wu, Wei Ju, Xiao Luo, and Ming Zhang. 2024. A Diffusion Model for POI Recommendation. ACM Trans. Inf. Syst. 42, 2 (2024), 54:1–54:27. https://doi.org/10.1145/3624475
- Rendle et al. (2009) Steffen Rendle, Christoph Freudenthaler, Zeno Gantner, and Lars Schmidt-Thieme. 2009. BPR: Bayesian Personalized Ranking from Implicit Feedback. In UAI 2009, Proceedings of the Twenty-Fifth Conference on Uncertainty in Artificial Intelligence, Montreal, QC, Canada, June 18-21, 2009, Jeff A. Bilmes and Andrew Y. Ng (Eds.). AUAI Press, 452–461. https://www.auai.org/uai2009/papers/UAI2009_0139_48141db02b9f0b02bc7158819ebfa2c7.pdf
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-Resolution Image Synthesis with Latent Diffusion Models. In IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2022, New Orleans, LA, USA, June 18-24, 2022. IEEE, 10674–10685. https://doi.org/10.1109/CVPR52688.2022.01042
- Sarwar et al. (2001) Badrul Munir Sarwar, George Karypis, Joseph A. Konstan, and John Riedl. 2001. Item-based collaborative filtering recommendation algorithms. In Proceedings of the Tenth International World Wide Web Conference, WWW 10, Hong Kong, China, May 1-5, 2001, Vincent Y. Shen, Nobuo Saito, Michael R. Lyu, and Mary Ellen Zurko (Eds.). ACM, 285–295. https://doi.org/10.1145/371920.372071
- Shen et al. (2024) Xudong Shen, Chao Du, Tianyu Pang, Min Lin, Yongkang Wong, and Mohan S. Kankanhalli. 2024. Finetuning Text-to-Image Diffusion Models for Fairness. In The Twelfth International Conference on Learning Representations, ICLR 2024, Vienna, Austria, May 7-11, 2024. OpenReview.net. https://openreview.net/forum?id=hnrB5YHoYu
- Shenbin et al. (2020) Ilya Shenbin, Anton Alekseev, Elena Tutubalina, Valentin Malykh, and Sergey I. Nikolenko. 2020. RecVAE: A New Variational Autoencoder for Top-N Recommendations with Implicit Feedback. In WSDM ’20: The Thirteenth ACM International Conference on Web Search and Data Mining, Houston, TX, USA, February 3-7, 2020, James Caverlee, Xia (Ben) Hu, Mounia Lalmas, and Wei Wang (Eds.). ACM, 528–536. https://doi.org/10.1145/3336191.3371831
- Steck (2019) Harald Steck. 2019. Embarrassingly Shallow Autoencoders for Sparse Data. In The World Wide Web Conference, WWW 2019, San Francisco, CA, USA, May 13-17, 2019, Ling Liu, Ryen W. White, Amin Mantrach, Fabrizio Silvestri, Julian J. McAuley, Ricardo Baeza-Yates, and Leila Zia (Eds.). ACM, 3251–3257. https://doi.org/10.1145/3308558.3313710
- Sun et al. (2020) Jianing Sun, Wei Guo, Dengcheng Zhang, Yingxue Zhang, Florence Regol, Yaochen Hu, Huifeng Guo, Ruiming Tang, Han Yuan, Xiuqiang He, and Mark Coates. 2020. A Framework for Recommending Accurate and Diverse Items Using Bayesian Graph Convolutional Neural Networks. In KDD ’20: The 26th ACM SIGKDD Conference on Knowledge Discovery and Data Mining, Virtual Event, CA, USA, August 23-27, 2020, Rajesh Gupta, Yan Liu, Jiliang Tang, and B. Aditya Prakash (Eds.). ACM, 2030–2039. https://doi.org/10.1145/3394486.3403254
- Togashi et al. (2024) Riku Togashi, Kenshi Abe, and Yuta Saito. 2024. Scalable and Provably Fair Exposure Control for Large-Scale Recommender Systems. In Proceedings of the ACM on Web Conference 2024, WWW 2024, Singapore, May 13-17, 2024, Tat-Seng Chua, Chong-Wah Ngo, Ravi Kumar, Hady W. Lauw, and Roy Ka-Wei Lee (Eds.). ACM, 3307–3318. https://doi.org/10.1145/3589334.3645390
- Vassøy and Langseth (2024) Bjørnar Vassøy and Helge Langseth. 2024. Consumer-side fairness in recommender systems: a systematic survey of methods and evaluation. Artif. Intell. Rev. 57, 4 (2024), 101. https://doi.org/10.1007/S10462-023-10663-5
- Vignac et al. (2023) Clément Vignac, Igor Krawczuk, Antoine Siraudin, Bohan Wang, Volkan Cevher, and Pascal Frossard. 2023. DiGress: Discrete Denoising diffusion for graph generation. In The Eleventh International Conference on Learning Representations, ICLR 2023, Kigali, Rwanda, May 1-5, 2023. OpenReview.net. https://openreview.net/forum?id=UaAD-Nu86WX
- Walker et al. (2022) Joojo Walker, Ting Zhong, Fengli Zhang, Qiang Gao, and Fan Zhou. 2022. Recommendation via Collaborative Diffusion Generative Model. In Knowledge Science, Engineering and Management - 15th International Conference, KSEM 2022, Singapore, August 6-8, 2022, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 13370), Gérard Memmi, Baijian Yang, Linghe Kong, Tianwei Zhang, and Meikang Qiu (Eds.). Springer, 593–605. https://doi.org/10.1007/978-3-031-10989-8_47
- Wang et al. (2017) Jun Wang, Lantao Yu, Weinan Zhang, Yu Gong, Yinghui Xu, Benyou Wang, Peng Zhang, and Dell Zhang. 2017. IRGAN: A Minimax Game for Unifying Generative and Discriminative Information Retrieval Models. In Proceedings of the 40th International ACM SIGIR Conference on Research and Development in Information Retrieval, Shinjuku, Tokyo, Japan, August 7-11, 2017, Noriko Kando, Tetsuya Sakai, Hideo Joho, Hang Li, Arjen P. de Vries, and Ryen W. White (Eds.). ACM, 515–524. https://doi.org/10.1145/3077136.3080786
- Wang et al. (2023) Wenjie Wang, Yiyan Xu, Fuli Feng, Xinyu Lin, Xiangnan He, and Tat-Seng Chua. 2023. Diffusion Recommender Model. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2023, Taipei, Taiwan, July 23-27, 2023, Hsin-Hsi Chen, Wei-Jou (Edward) Duh, Hen-Hsen Huang, Makoto P. Kato, Josiane Mothe, and Barbara Poblete (Eds.). ACM, 832–841. https://doi.org/10.1145/3539618.3591663
- Wu et al. (2024) Likang Wu, Zhi Zheng, Zhaopeng Qiu, Hao Wang, Hongchao Gu, Tingjia Shen, Chuan Qin, Chen Zhu, Hengshu Zhu, Qi Liu, Hui Xiong, and Enhong Chen. 2024. A survey on large language models for recommendation. World Wide Web (WWW) 27, 5 (2024), 60. https://doi.org/10.1007/S11280-024-01291-2
- Wu et al. (2023) Shiwen Wu, Fei Sun, Wentao Zhang, Xu Xie, and Bin Cui. 2023. Graph Neural Networks in Recommender Systems: A Survey. ACM Comput. Surv. 55, 5 (2023), 97:1–97:37. https://doi.org/10.1145/3535101
- Xu et al. (2022) Minkai Xu, Lantao Yu, Yang Song, Chence Shi, Stefano Ermon, and Jian Tang. 2022. GeoDiff: A Geometric Diffusion Model for Molecular Conformation Generation. In The Tenth International Conference on Learning Representations, ICLR 2022, Virtual Event, April 25-29, 2022. OpenReview.net. https://openreview.net/forum?id=PzcvxEMzvQC
- Yang et al. (2016) Dingqi Yang, Daqing Zhang, and Bingqing Qu. 2016. Participatory Cultural Mapping Based on Collective Behavior Data in Location-Based Social Networks. ACM Trans. Intell. Syst. Technol. 7, 3 (2016), 30:1–30:23. https://doi.org/10.1145/2814575
- Yang et al. (2024) Ling Yang, Zhilong Zhang, Yang Song, Shenda Hong, Runsheng Xu, Yue Zhao, Wentao Zhang, Bin Cui, and Ming-Hsuan Yang. 2024. Diffusion Models: A Comprehensive Survey of Methods and Applications. ACM Comput. Surv. 56, 4 (2024), 105:1–105:39. https://doi.org/10.1145/3626235
- Yu et al. (2024) Junliang Yu, Xin Xia, Tong Chen, Lizhen Cui, Nguyen Quoc Viet Hung, and Hongzhi Yin. 2024. XSimGCL: Towards Extremely Simple Graph Contrastive Learning for Recommendation. IEEE Trans. Knowl. Data Eng. 36, 2 (2024), 913–926. https://doi.org/10.1109/TKDE.2023.3288135
- Zhao et al. (2022b) Minghao Zhao, Le Wu, Yile Liang, Lei Chen, Jian Zhang, Qilin Deng, Kai Wang, Xudong Shen, Tangjie Lv, and Runze Wu. 2022b. Investigating Accuracy-Novelty Performance for Graph-based Collaborative Filtering. In SIGIR ’22: The 45th International ACM SIGIR Conference on Research and Development in Information Retrieval, Madrid, Spain, July 11 - 15, 2022, Enrique Amigó, Pablo Castells, Julio Gonzalo, Ben Carterette, J. Shane Culpepper, and Gabriella Kazai (Eds.). ACM, 50–59. https://doi.org/10.1145/3477495.3532005
- Zhao et al. (2022a) Wayne Xin Zhao, Yupeng Hou, Xingyu Pan, Chen Yang, Zeyu Zhang, Zihan Lin, Jingsen Zhang, Shuqing Bian, Jiakai Tang, Wenqi Sun, Yushuo Chen, Lanling Xu, Gaowei Zhang, Zhen Tian, Changxin Tian, Shanlei Mu, Xinyan Fan, Xu Chen, and Ji-Rong Wen. 2022a. RecBole 2.0: Towards a More Up-to-Date Recommendation Library. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management, Atlanta, GA, USA, October 17-21, 2022, Mohammad Al Hasan and Li Xiong (Eds.). ACM, 4722–4726. https://doi.org/10.1145/3511808.3557680